A programmer plans to develop a new software system. In planning for the operating system that he will use, he needs to estimate the percentage of computers that use a new operating system. How many computers must be surveyed in order to be 99% confident that his estimate is in error by no more than five percentage points? Complete parts (a) through (c) below. a) Assume that nothing is known about the percentage of computers with new operating systems (Round up to the nearest integer.)
Answer #1

Solution:

Given that,

\hat p\hatp = 0.5

1 - \hat p = 1 - 0.5 = 0.5

margin of error = E = 5% = 0.05

At 99% confidence level the z is ,

\alpha = 1 - 99% = 1 - 0.99 = 0.01

\alpha / 2 = 0.01 / 2 = 0.005

Z\alpha/2 = Z0.005 = 2.576

Sample size = n = ((Z\alpha / 2) / E)2 * \hat p * (1 - \hat p)

= (2.576 / 0.05)2 * 0.5 * 0.5

= 663.5776

= 664

n = sample size = 664